OpenStudy (anonymous):

The monthly cost of driving a car depends on the number of miles driven. Lynn found that in May her driving cost was $380 for 580 mi and in June her cost was $460 for 900 mi. Assume that there is a linear relationship between the monthly cost C of driving a car and the distance driven d. (a) Find a linear equation that relates C and d. (b) Use part (a) to predict the cost of driving 1700 mi per month. (c) Draw the graph of the linear equation.

OpenStudy (anonymous):

alright which do you think would be the dependent variable?

OpenStudy (anonymous):

c or d?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

alright so that'll be our "y" and d will be our "x"

OpenStudy (anonymous):

(580,380) and (900, 460)

OpenStudy (anonymous):

Can you find the slope?

OpenStudy (anonymous):

.25

OpenStudy (anonymous):

Alright, same as 1/4. Now using the slope and one of the points we can write slope point form. which point do you want to use?

OpenStudy (anonymous):

Still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(580,380)

OpenStudy (anonymous):

Alright. are you familiar with point slope form?

OpenStudy (anonymous):

\[y-y _{_{0}}=m(x-x _{0})\]

OpenStudy (anonymous):

In our case we use c for y and d for x

OpenStudy (anonymous):

So we get c-380=1/4(d-580)

OpenStudy (anonymous):

Distribute the 1/4 and we get c-380=1/4d-145

OpenStudy (anonymous):

f(c)=4c-940

OpenStudy (anonymous):

Add 380 to both sides and we get: c=1/4d+235

OpenStudy (anonymous):

understand so far?

OpenStudy (anonymous):

yah so what is the final equation?

OpenStudy (anonymous):

a) Find a linear equation that relates C and d.

OpenStudy (anonymous):

c=1/4d+235 is the answer to a

OpenStudy (anonymous):

Ready to do b?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

wait i got part b :) what does the graph look like?

OpenStudy (anonymous):

Hmm, well the y intercept is (0,235)

OpenStudy (anonymous):

can u please graph it

OpenStudy (anonymous):